OpenStudy (anonymous):

A motorboat leaves harbor and travels at an average speed of 9mph toward a smaal island. Two hours later a cabin cruiser leaves the same harbor and travels at an average speed of 18mph toward the same island. In how many hours after the cabin cruiser leaves the harbor will it be alongside the motorboat?

OpenStudy (anonymous):

2hour

OpenStudy (anonymous):

How to get 2hours?

OpenStudy (anonymous):

total distance between them=18 miles

OpenStudy (anonymous):

and their relative speed=18-9=9

OpenStudy (anonymous):

Es

OpenStudy (anonymous):

Yes I mean

OpenStudy (anonymous):

so time =distance between them /speed

OpenStudy (anonymous):

18/9=2

OpenStudy (anonymous):

Thanks a lot

OpenStudy (anonymous):

welcome...)